Simon Company’s year-end balance sheets follow. At December 31 Current Yr 1 Yr Ago 2 Yrs Ago Assets Cash $ 28,921 $ 33,806 $ 35,215 Accounts receivable, net 89,500 62,500 51,800 Merchandise inventory 112,500 84,500 51,000 Prepaid expenses 9,314 8,874 3,913 Plant assets, net 254,909 237,168 217,372 Total assets $ 495,144 $ 426,848 $ 359,300 Liabilities and Equity Accounts payable $ 122,058 $ 71,416 $ 46,953 Long-term notes payable secured by mortgages on plant assets 92,156 97,193 78,611 Common stock, $10 par value 162,500 162,500 162,500 Retained earnings 118,430 95,739 71,236 Total liabilities and equity $ 495,144 $ 426,848 $ 359,300 The company’s income statements for the Current Year and 1 Year Ago, follow. Assume that all sales are on credit: For Year Ended December 31 Current Yr 1 Yr Ago Sales $ 643,687 $ 507,949 Cost of goods sold $ 392,649 $ 330,167 Other operating expenses 199,543 128,511 Interest expense 10,943 11,683 Income tax expense 8,368 7,619 Total costs and expenses 611,503 477,980 Net income $ 32,184 $ 29,969 Earnings per share $ 1.98 $ 1.84

(2-a) Compute accounts receivable turnover. (2-b) For each ratio, determine if it improved or worsened in the current year.
Answer #1
2a
Accounts Receivable Turnover
Choose Numerator: / Choose Denominator: = Accounts Receivable Turnover
Net sales / Average accounts receivable, net = Accounts receivable turnover
Current Yr 643687 / 76000 = 8.5 times
1 Yr Ago 507949 / 57150 = 8.9 times
2b
Accounts receivable turnover worsened in the current year
Workings:
Average accounts receivable, net
Current Yr 76000 =(89500+62500)/2
1 Yr Ago 57150 =(62500+51800)/2